-2/3x=8 this problem is a linear equation problem.From algebra 2

Answer by checkley71(8405)

(-2/3)X=8
-2X=24
X=24/-2
X=-12 SOLUTION OR
-2/(3X)=8
-2=24X
X=-2/24
X=-1/12 SOLUTION